Sales price = $60 Variable cost = $30 What is the contribution margin ratio?

Sales price = $60

Variable cost = $30

What is the contribution margin ratio?
Answer #1

Explanation

​​​​​​sale price 60

Variable cost 30

The contribution margin ratio=?

Contribution margin=net sales-variable cost

Contribution margin=60-30=30

Contribution margin ratio=contribution margin/sale price

Contribution margin ratio=30/60=0.5

In percentage 0.5*100=50%